# Weighted Birkhoff averages: Deterministic and probabilistic perspectives

### v2

@@ 1.1.3 Exponentially computing Fourier coefficients using weighted Birkhoff averages
$F\left({K\left(\theta\right)}\right)=K\left({\theta+\rho}\right),\;\;\text{where $\rho\in\mathbb{T}^{d}$ is a nonresonant rotation.}$ (1.7)
@@
Figure 5: The geometric intuition of the conjugation (1.7) (with $d=1$) and commutative diagram
@@
$K\left(x\right)={\left({{K^{1}}\left(x\right),\ldots,{K^{D}}\left(x\right)}\right)^{\top}}=\sum\limits_{k\in{\mathbb{Z}^{d}}}{\hat{K}\left(k\right)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}\;\;\text{or}\;\;\sum\limits_{k\in\mathbb{Z}_{*}^{\infty}}{\hat{K}\left(k\right)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}$
@@
$\hat{K}\left(s\right)=\int_{{\mathbb{T}^{d}}}{K\left(x\right)\exp\left({-2\pi i\left\langle{s,x}\right\rangle}\right)dx}.$
@@
is well-defined and equals $\hat{u}(0)$, see [Bou05, MP21, CGP24], for instance. Set ${\theta_{n}}:={\theta_{0}}+n\rho$ for $n\in\mathbb{N}$. Then, by induction, we obtain from the conjugation (1.7) that
@@
$\hat{K}\left(s\right)\approx\exp\left({-2\pi i\left\langle{s,{\theta_{0}}}\right\rangle}\right)\cdot\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right){p_{n}}\exp\left({-2\pi in\left\langle{s,\rho}\right\rangle}\right)},\;\;N\gg 1,$ (1.9)
@@
${\left\|{\hat{K}\left(s\right)-\exp\left({-2\pi i\left\langle{s,{\theta_{0}}}\right\rangle}\right)\cdot\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right){p_{n}}\exp\left({-2\pi in\left\langle{s,\rho}\right\rangle}\right)}}\right\|_{{\ell^{\infty}}}}=o\left(1\right)$ (1.10)
@@ 1.2.1 Weighted laws of large numbers and weighted strong laws of large numbers
    $\mathop{\lim}\limits_{N\to+\infty}{\bf P}\left({\left|{\frac{1}{{{A_{N}}}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){X_{n}}}-\frac{1}{{{A_{N}}}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){\bf E}{X_{n}}}}\right|<\varepsilon}\right)=1.$
@@
    $\mathop{\lim}\limits_{N\to+\infty}{\bf P}\left({\left|{\frac{1}{{{A_{N}}}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){X_{n}}}-\frac{1}{{{A_{N}}}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){\bf E}{X_{n}}}}\right|<\varepsilon}\right)=1.$
Secondly, it should be emphasized that the following weighted strong laws of large numbers (Theorems 1.6 and 1.7 below) and the weighted central limit theorem (Theorem 1.8 in the next section), are primarily built upon historical results [Thr87, BC00, CG07, Kla09] and have the potential for further refinement.
@@ 1.2.2 Weighted central limit theorem
which is precisely the upper bound order we have obtained in Theorem 1.8. However, it is worth emphasizing that the weighting coefficients that can accelerate the rate of the weighted central limit theorem to $\mathcal{O}\left(N^{-1}\right)$ may not necessarily result in an exponential acceleration in the deterministic case, for example, when taking ${\theta_{n}}=\sin^{2}\left({\pi n/N}\right)/\sum\nolimits_{j=1}^{N}{\sin^{2}\left({\pi j/N}\right)}$ for $1\leqslant n\leqslant N$.
This weighted approach can also be extended to establish a variety of additional results, which we recommend exploring further in [MM07, Corollary 4], [GS09, Theorem 2.1], [KS12, Theorem 1.2], and also [Gru14, Theorem 1.1]. From a probabilistic perspective, under the assumption that the fourth moment is finite (but without the additional requirements in Theorem 1.8), the weights that achieve the weighted central limit theorem at a rate of $\mathcal{O}(N^{-1})$ are of almost full normalized Lebesgue measure $\mathfrak{s}_{N-1}$ on $S^{N-1}$ (except for a set of small measure), see [KS12, Theorem 1.1]; while under the assumption that the fifth moment is finite, by introducing a corrected normal distribution, it can be shown that the weights that achieve the weighted central limit theorem at a rate of $\mathcal{O}(N^{-3/2})$ are of almost full normalized Lebesgue measure $\mathfrak{s}_{N-1}$ on $S^{N-1}$, see [Bob20, Theorem 5.1.1]. Unfortunately, we do not know whether our weight derived from (1.1) or some other variants (e.g., positive on $[0,1/2]$ and negative on $[1/2,1]$, see more general cases in [TL24b]) with the $C_{0}^{\infty}\left([0,1]\right)$ property as well as the normalization property belongs to such almost full measure sets. One feasible approach is to verify certain arithmetic properties of the weights, such as those described in [KS12, (iii)] (or the assumptions within Conjecture 1.1 below). This approach requires further exploration and is of significant interest. Here, based on [KS12, Theorem 1.1] in the measure sense, we leave the following conjecture:
@@ 2 New contributions and historical comparisons
Theorem 1.2 was conceived with two motivations: first, to summarize the existing results, and second, to establish improved quantitative versions, both of which lend themselves to practical simulations in dynamical systems. More importantly, Theorem 1.2 allows us to investigate, at a theoretical level, how the convergence of weighted Birkhoff averages is shaped by the interplay among the weighting function, nonresonance, and regularity. For these reasons, we prefer to regard the bulk of Theorem 1.2 as an essential contribution of this paper. Specifically, Theorem 1.2 treats the general weighting function $w_{p,q}$, going beyond the special case $w_{1,1}$ that has dominated earlier theoretical works [DSSY17, DY18, DM23, TL24a, TL24b, TL24c]. This provides rigorous, quantitative guarantees for the numerical simulations explored in [DSSY17, CCGd24], which are of considerable practical importance******It is worth emphasizing that the theoretical analysis becomes far more intricate than the $p=q=1$ case, and the techniques previously employed are no longer effective.. A detailed comparison with the historical literature will be given in the first paragraph of Section 3.2; here we confine ourselves to a brief summary. Indeed, the polynomial type estimate (I)–(1) is an immediate consequence of the techniques developed in [DSSY17, DY18, DM23, TL24a], requiring no essential alteration, whereas the exponential type estimate (I)–(3) was established in [TL25a]. Hence, neither forms part of the original contributions of this paper. The exponential type results (I)–(2), (I)–(4), (II)–(5), (II)–(6), and (III)–(7) are new. We emphasize, however, that for the special case $p=q=1$ the statement corresponding to (I)–(2) was obtained in [TL25a]; therefore we prefer not to regard our treatment of general $p,q>0$ as an essential innovation. For the special case $p=q=1$, a qualitative version of (II)–(5) was established in [TL24c] (also in a multiple sense); here we improve it quantitatively for general $p,q>0$, a highly nontrivial yet fundamental challenge that constitutes an essential novelty. Similarly, for $p=q=1$, a weaker, qualitative form of (II)–(6) was given in [TL24a, TL24c]. Finally, (III)–(7) represents another essential novelty: we provide the first quantitative exponential convergence result for the purely periodic case. To further illustrate the contrast between our exponential results and earlier ones, consider the most commonly used quasi-periodic case (I)–(2). When $p=q=1$, [TL25a] proves that the weighted Birkhoff average converges at the rate $\mathcal{O}\left(\exp(-N^{\zeta})\right)$ for any fixed $0<\zeta<(d+2)^{-1}$. Our work shows that for general $p,q>0$ the exponent $\zeta$ depends explicitly on these parameters via any fixed $0<\zeta<(d+1+1/\min\{p,q\})^{-1}$, a refinement of both theoretical and practical importance. Similarly, in the most important almost-periodic setting (II)–(6), the qualitative result of [TL24c] guarantees the weighted Birkhoff average to converge at the rate $\mathcal{O}\left(\exp\left(-(\log N)^{\zeta}\right)\right)$ for $p=q=1$, yet $\zeta$ is only known to be positive, with no explicit connection to the underlying infinite-dimensional spatial structure. Our new result (II)–(6) makes this dependence transparent as $2\leqslant{\zeta}<1+\eta$††††††Here $2\leqslant\eta\in\mathbb{N}^{+}$ characterizes the properties of the infinite-dimensional spatial structure. and is therefore of significance in infinite-dimensional ergodic theory.
@@
Theorem 1.3, although not regarded as the essential contribution of the present paper–we defer its complete proof to a companion work–is nevertheless a novel result. Weighted Birkhoff averages were originally introduced to accelerate the convergence of the classical Birkhoff average, both theoretically and in practical simulations. In all preceding studies (as cited in the Introduction), the weighting functions appearing in Theorem 1.3 and Remark 1.3 have never been considered, yet they possess markedly stronger theoretical acceleration properties than those currently in use. This reveals the profound link between the asymptotics of the weighting function and the convergence rate of the weighted Birkhoff average, thereby lending the result additional importance.
@@
Theorem 1.4 is entirely new and constitutes one of the essential contributions of this paper. It is worth stressing that this result is intimately tied to the paper’s overarching focus on convergence. Weighted Birkhoff averages have proved remarkably effective in concrete computations–notably for calculations of Fourier coefficients, as surveyed in Section 1.1.2–yet no rigorous theory has hitherto guaranteed their correctness, let alone established an exponential convergence result. Moreover, existing applications are confined to the quasi-periodic setting; the almost-periodic case has remained untouched. Motivated by these gaps, we establish an almost complete, quantitative theorem that is simultaneously amenable to practical use. A key innovation is the introduction of the notion “effective order”, which quantifies, for the first time, the range of modes whose Fourier coefficients can be calculated with exponential convergence via weighted Birkhoff averages. This aspect has received virtually no prior consideration. Finally, we emphasize that Theorem 1.4 is not a mere corollary of Theorem 1.2. In fact, parts of the proof of the latter are subsumed within the more intricate argument required for the former, which we have explained in Remark 1.4.

### v3

@@ 1.1.3 Exponentially computing Fourier coefficients using weighted Birkhoff averages
$F\left({K\left(\theta\right)}\right)=K\left({\theta+\rho}\right),\quad\text{where $\rho\in\mathbb{T}^{d}$ is a nonresonant rotation.}$ (1.7)
@@
Figure 5: The geometric intuition of the conjugation (1.7) (with $d=1$) and the commutative diagram
@@
$K\left(x\right)={\left({{K^{1}}\left(x\right),\ldots,{K^{D}}\left(x\right)}\right)^{\top}}=\sum\limits_{k\in{\mathbb{Z}^{d}}}{\widehat{K}\left(k\right)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}\quad\text{or}\quad\sum\limits_{k\in\mathbb{Z}_{*}^{\infty}}{\widehat{K}\left(k\right)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}$
@@
$\widehat{K}\left(s\right)=\int_{{\mathbb{T}^{d}}}{K\left(x\right)\exp\left({-2\pi i\left\langle{s,x}\right\rangle}\right)dx}.$
@@
is well-defined and equals $\widehat{u}(0)$, see [Bou05, MP21, CGP24], for instance. Set ${\theta_{n}}:={\theta_{0}}+n\rho$ for $n\in\mathbb{N}$. Then, by induction, we obtain from the conjugation (1.7) that
@@
$\widehat{K}\left(s\right)\approx\exp\left({-2\pi i\left\langle{s,{\theta_{0}}}\right\rangle}\right)\cdot\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right){p_{n}}\exp\left({-2\pi in\left\langle{s,\rho}\right\rangle}\right)},\quad N\gg 1,$ (1.9)
@@
${\left\|{\widehat{K}\left(s\right)-\exp\left({-2\pi i\left\langle{s,{\theta_{0}}}\right\rangle}\right)\cdot\frac{1}{{A_{N}^{p,q}}}\sum\limits_{n=1}^{N-1}{{w_{p,q}}\left({n/N}\right){p_{n}}\exp\left({-2\pi in\left\langle{s,\rho}\right\rangle}\right)}}\right\|_{{\ell^{\infty}}}}=o\left(1\right)$ (1.10)
@@
1.1.4 Which factors may affect the “practical” convergence rate?
In contrast to the dynamical-systems perspective, researchers focusing on numerical simulations are typically more interested in the “practical” convergence rate for finite yet large $N$. This motivates the following question: Which factors may affect the “practical” convergence rate of the weighted quasi-Monte Carlo method investigated in this paper? While [DSSY17, Sections 4 and 5] offer relevant insights, we provide in this section a deeper and more theoretical discussion.
Let us consider the weighted Birkhoff average in (1.2) utilizing the weighted quasi-Monte Carlo method with $p,q>0$. As in Theorem 1.2, we continue to examine its uniform convergence rate to $\int_{\mathbb{T}^{d}}fdx$ in the $\ell^{\infty}$-norm. In order to clarify the presentation, we focus exclusively on the case $d\in\mathbb{N}^{+}$, and our convergence analysis is restricted to the polynomial type of finite order§§§Given that this paper aims to make this weighted quasi-Monte Carlo method more user-friendly, a more precise quantitative analysis of exponential convergence (from a more dynamical-systems perspective) will be provided as a byproduct in a separate paper focusing on large deviations.. According to [DSSY17, DY18, DM23, TL24a], the convergence rate is $\mathcal{O}\left(N^{-m}\right)$ when the regularity of the observable $f$ and the nonresonance of the rotation $\rho$ satisfies
$\sum\limits_{0\neq k\in{\mathbb{Z}^{d}}}{{{\|{\widehat{f}(k)}\|}_{{\ell^{\infty}}}}\left({\mathop{\sup}\limits_{n\in\mathbb{Z}}\left({{{\left|{\left\langle{k,\rho}\right\rangle-n}\right|}^{-m}}}\right)}\right)}<+\infty,\quad 2\leqslant m\in\mathbb{N}^{+}.$
Observe that if $\rho$ is nonresonant, we can assume without loss of generality that it satisfies the following nonresonance condition
$\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\frac{\gamma}{{\Delta\left({{{\left\|k\right\|}_{{\ell^{1}}}}}\right)}},\quad\gamma>0,\quad\forall 0\neq k\in{\mathbb{Z}^{d}},\quad\forall n\in\mathbb{Z},$
where $\Delta$ is an approximating function (see Remark 1.3) with $\Delta(1)>0$. In this case, using a slightly modified argument in [DSSY17, DY18] (for instance, [TL24a, Section 4.1]), the convergence rate can be estimated as $\big({\prod\nolimits_{j=1}^{4}{{{\mathcal{C}}_{j}}}}\big)N^{-m}$ ($2\leqslant m\in\mathbb{N}^{+}$), where the coefficients are explicitly given by
${\mathcal{C}_{1}}=\mathop{\sup}\limits_{N\in{\mathbb{N}^{+}}}\frac{N}{{A_{N}^{p,q}}},\quad{\mathcal{C}_{2}}={\left\|{{D^{m}}{w_{p,q}}}\right\|_{{L^{1}}\left({0,1}\right)}},\quad{\mathcal{C}_{3}}=\frac{1}{{{{\left({2\pi\gamma}\right)}^{m}}}},\quad{\mathcal{C}_{4}}=\sum\limits_{0\neq k\in{\mathbb{Z}^{d}}}{{{\|{\widehat{f}(k)}\|}_{{\ell^{\infty}}}}{\Delta^{m}}\left({{{\left\|k\right\|}_{{\ell^{1}}}}}\right)}.$
At this point, we can identify which factors may affect the convergence rate. Please keep in mind that this serves only as an intuitive understanding and a possible explanation for numerical simulations, rather than a rigorous mathematical proof—because the convergence rate discussed here is not necessarily the actual rate, but merely an upper bound. To study this problem rigorously, one must simultaneously consider the lower bound of the convergence rate, which requires entirely different techniques and is considerably more challenging. This will be detailed in [TL25c]. However, from the perspective of numerical simulations, the following interpretation from a dynamical-systems perspective may be helpful. Indeed, the following analysis has been empirically validated as correct in practice.
  • (I)
    In numerical simulations, the convergence rate is sometimes approximated by finite-order polynomials such as $N^{-m}$ (even though it might actually be exponential) for finite yet large $N$. Hence, the specific value of $m$ is a potential factor.
  • (II)
    The coefficients $\mathcal{C}_{1}$} and $\mathcal{C}_{2}$} reveal that the weighting function $w_{p,q}(x)$ itself is a potential factor. Straightforward verification shows $\mathop{\lim}\nolimits_{N\to+\infty}N/A_{N}^{p,q}=1$ (see Section 3.3), but for large $p$ and $q$, finite $N$ may cause $\mathcal{C}_{1}$} to deviate significantly from $1$. As for $\mathcal{C}_{2}$}, [TL25b, Lemma 4.1] provides its asymptotic estimate with respect to $m$, allowing for a clearer assessment of its potential impact.
  • (III)
    The coefficients $\mathcal{C}_{3}$} and $\mathcal{C}_{4}$} reveal that the rotation $\rho$ is also a potential factor. This can be explained from two aspects. First, for a given nonresonance condition—that is, with $\Delta$ fixed—a smaller $\gamma$ leads to a larger $\mathcal{C}_{3}$}. Second, for a fixed $\gamma$, the type of nonresonance—such as Diophantine, Brjuno, or Liouville—also affects $\mathcal{C}_{4}$}. In particular, if we further consider a given observable $f$, then from $\mathcal{C}_{4}$} we can intuitively see that weaker nonresonance tends to slow down the “practical” convergence rate. For example, when $\rho$ is of Diophantine type, i.e., $\Delta(x)=x^{\tau}$ with some $\tau\geqslant d$, then $C_{4}$ becomes $\sum\nolimits_{0\neq k\in{\mathbb{Z}^{d}}}{{{\|{\widehat{f}(k)}\|}_{{\ell^{\infty}}}}\left\|k\right\|_{{\ell^{1}}}^{\tau}}$. Clearly, sufficiently fast polynomial decay of the Fourier coefficients of $f$ ensures that $C_{4}<+\infty$. However, if $\rho$ is of Liouville type, such as when $\Delta(x)=\exp(x)$ (and there simultaneously exist infinitely many tuples $\left({k,n}\right)\in{\mathbb{Z}^{d}}\times\mathbb{Z}$ such that $\left|{\left\langle{k,\rho}\right\rangle-n}\right|\sim\gamma\exp\left({-{{\left\|k\right\|}_{{\ell^{1}}}}}\right)$), then to guarantee $C_{4}=\sum\nolimits_{0\neq k\in{\mathbb{Z}^{d}}}{{{\|{\widehat{f}(k)}\|}_{{\ell^{\infty}}}}\exp\left({{{\left\|k\right\|}_{{\ell^{1}}}}}\right)}<+\infty$, we have to require exponential decay of the Fourier coefficients of $f$. Theoretical results concerning Liouville nonresonance have been extensively discussed in [TL24a, TL24b, TL25a]. [DSSY17] also corroborates these viewpoints through numerical simulations. As a remark, in practice there appears to be no feasible way to decide whether $\rho$ is Liouville; one usually has to rely on empirical judgementZ. Tong first came across this point during a 2024 talk by Prof. Alessandra Celletti related to weighted Birkhoff averages.. Consequently, both aspects above are indispensable. For instance, even if $\rho$ is Diophantine, an extremely small $\gamma$ may still tempt one to place it, intuitively, in the Liouville class—although, from a dynamical-systems perspective, the two situations are entirely different.
  • (IV)
    The coefficient $\mathcal{C}_{4}$} reveals that $f$ is also a potential factor. This can be examined from two perspectives. On the one hand, given a fixed rotation $\rho$, it is clear that higher regularity of $f$ increases the possibility of $\mathcal{C}_{4}$$\infty$. However, this may not significantly affect the size of $\mathcal{C}_{4}$}, since regularity primarily affects the decay of Fourier coefficients at high frequencies, while the value of $\mathcal{C}_{4}$} is partly determined by low-frequency Fourier coefficients. On the other hand, the only rigorous statement that can be made here concerns how the “scale” of $f$ affects the “practical” convergence rate—for example, replacing $f$ with $2f$. The homogeneity of the weighted quasi-Monte Carlo method with respect to $f$ provides a direct explanation in such cases.
In conclusion, the “practical” convergence rate may be affected by the specific value of $m$, as well as all dynamical system-related factors involved in this weighted quasi-Monte Carlo method—namely, the weighting function, the rotation, and the observable. It can be seen that the most critical factors are the nonresonance of the rotation and the regularity of the observable, which precisely correspond to the central perspectives of interest in dynamical systems.
@@ 1.2.1 Weighted laws of large numbers and weighted strong laws of large numbers
The weighted laws of large numbers and the weighted strong laws of large numbers have long been central topics in probability theory. A substantial body of historical and recent advances exists on these topics; see, for instance, [JOP65, Pru66, Roh71, Thr87, BC00, Jaj03, CG07, JL08, LM10, CS16, FMZ17, NMSA20, Ad24] and the references therein.
@@
    $\mathop{\lim}\limits_{N\to+\infty}{\bf P}\left({\left|{\frac{1}{{{{A_{N}^{p,q}}}}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){X_{n}}}-\frac{1}{A_{N}^{p,q}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){\bf E}{X_{n}}}}\right|<\varepsilon}\right)=1.$
@@
    $\mathop{\lim}\limits_{N\to+\infty}{\bf P}\left({\left|{\frac{1}{A_{N}^{p,q}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){X_{n}}}-\frac{1}{A_{N}^{p,q}}\sum\limits_{n=1}^{N}{{w_{p,q}}\left({n/N}\right){\bf E}{X_{n}}}}\right|<\varepsilon}\right)=1.$
Secondly, it should be emphasized that the following weighted strong laws of large numbers (Theorems 1.6 and 1.7 below) and the weighted central limit theorem (Theorem 1.8 in the next section), are primarily built upon historical results [Thr87, BC00, CG07, Kla09] and have the potential for further refinement—for instance, in conjunction with the previously mentioned works. In particular, these results can be further explored beyond the independent setting, allowing for dependent stochastic processes that will cover a broad range of nonlinear phenomena [FMZ17].
@@ 1.2.2 Weighted central limit theorem
Remark 1.5.
As previously mentioned, one may further consider the case of dependent stochastic processes, which can better capture nonlinear phenomena.
@@
which is precisely the upper bound order we have obtained in Theorem 1.8. However, it is worth emphasizing that the weighting coefficients that can accelerate the rate of the weighted central limit theorem to $\mathcal{O}\left(N^{-1}\right)$ may not necessarily result in an exponential acceleration in the deterministic case, for example, when taking ${\theta_{n}}=\sin^{2}\left({\pi n/N}\right)/\sum\nolimits_{j=1}^{N}{\sin^{2}\left({\pi j/N}\right)}$ for $1\leqslant n\leqslant N$ [TL25a].
This weighted approach can also be extended to establish a variety of additional results, which we recommend exploring further in [MM07, Corollary 4], [GS09, Theorem 2.1], [KS12, Theorem 1.2], and also [Gru14, Theorem 1.1]. From a probabilistic perspective, under the assumption that the fourth moment is finite (but without the additional requirements in Theorem 1.8), the weights that achieve the weighted central limit theorem at a rate of $\mathcal{O}(N^{-1})$ are of almost full normalized Lebesgue measure $\mathfrak{s}_{N-1}$ on $S^{N-1}$ (except for a set of small measure), see [KS12, Theorem 1.1]; while under the assumption that the fifth moment is finite, by introducing a corrected normal distribution, it can be shown that the weights that achieve the weighted central limit theorem at a rate of $\mathcal{O}(N^{-3/2})$ are of almost full normalized Lebesgue measure $\mathfrak{s}_{N-1}$ on $S^{N-1}$, see [Bob20, Theorem 5.1.1]. Unfortunately, we do not know whether our weights derived from (1.1) or some other variants (e.g., positive on $[0,1/2]$ and negative on $[1/2,1]$, see more general cases in [TL25a]) with the $C_{0}^{\infty}\left([0,1]\right)$ property as well as the normalization property belongs to such almost full measure sets. One feasible approach is to verify certain arithmetic properties of the weights, such as those described in [KS12, (iii)] (or the assumptions within Conjecture 1.1 below). This approach requires further exploration and is of significant interest. Here, based on [KS12, Theorem 1.1] in the measure sense, we leave the following conjecture:
@@ 2 New contributions and historical comparisons
[…]
@@
Theorem 1.3, although not regarded as the essential contribution of the present paper—we defer its complete proof to a companion work—is nevertheless a novel result. Weighted Birkhoff averages were originally introduced to accelerate the convergence of the classical Birkhoff average, both theoretically and in practical simulations. In all preceding studies (as cited in the Introduction), the weighting functions appearing in Theorem 1.3 and Remark 1.3 have never been considered, yet they possess markedly stronger theoretical acceleration properties than those currently in use. This reveals the profound link between the asymptotics of the weighting function and the convergence rate of the weighted Birkhoff average, thereby lending the result additional importance.
@@
Theorem 1.4 is entirely new and constitutes one of the essential contributions of this paper. It is worth stressing that this result is intimately tied to the paper’s overarching focus on convergence. Weighted Birkhoff averages have proved remarkably effective in concrete computations—notably for calculations of Fourier coefficients, as surveyed in Section 1.1.2—yet no rigorous theory has hitherto guaranteed their correctness, let alone established an exponential convergence result. Moreover, existing applications are confined to the quasi-periodic setting; the almost-periodic case has remained untouched. Motivated by these gaps, we establish an almost complete, quantitative theorem that is simultaneously amenable to practical use. A key innovation is the introduction of the notion “effective order”, which quantifies, for the first time, the range of modes whose Fourier coefficients can be calculated with exponential convergence via weighted Birkhoff averages. This aspect has received virtually no prior consideration. Finally, we emphasize that Theorem 1.4 is not a mere corollary of Theorem 1.2. In fact, parts of the proof of the latter are subsumed within the more intricate argument required for the former, which we have explained in Remark 1.4.